OpenStudy (anonymous):

determine if the proportion is true 6/12=10/18

OpenStudy (mattfeury):

find a common denominator and compare. in this case, 6 is a good denominator \[\frac{6}{12} = \frac{3}{6}\] \[\frac{10}{18} = \frac{10/6}{3}\] they are not equal so it is not true.

OpenStudy (mattfeury):

you could also try and "cross multiply," i.e. 6*18 = 12*10, which is also false.

OpenStudy (mattfeury):

thanks sandy :)

OpenStudy (anonymous):

6/12 is 1/2. The numerator of the right side would have to be 9 in order for the equation to be true.